Suppose that a 5.0£10¡5 M solution of a compound had a %T equal to 70.80 when measured in a
cuvette with a 1.0 cm path length. Find themolar extinction coefficient of this compound.
1
Expert's answer
2020-09-21T06:57:33-0400

The following expression shows the relationship between the absorbance and percent transmittance;

"A=2-log(\%T)"

Percent transmittance "(\%T)" "= 70.80"

Plug in the value into the formula

"A=2-log(70.80)\newline A=2-1.85=0.15"

Beer Lambert Law states that the amount of energy absorbed or transmitted by a solvent is directly proportional to the concentration of the solute and molar absorptivity of the solution i.e

"A=\epsilon lc \ where,"

"A=Absorbance=0.15\newline \epsilon=molar\ extinction\ co efficient\newline l=path \ length=1cm\newline C= concentration=5.0*10^{5} M" Rearrange the formula and make "\epsilon" the subject.

"\epsilon=\dfrac{A}{l*c}"

Substituting the values into the formula;

"\epsilon=\dfrac{0.15}{5*10^5M*1cm}=3000M^{-1}cm^{-1}"


Therefore "\epsilon""=3000M^{-1}cm^{-1}"